### v3

# Dynamics and topology of the Hat family of tilings

Abstract.
The recently discovered Hat tiling [17] admits a $4$-dimensional family of shape deformations, including the $1$-parameter family already known to yield alternate monotiles. The continuous hulls resulting from these tilings are all topologically conjugate dynamical systems, and hence have the same dynamics and topology. We construct and analyze a self-similar element of this family called the CAP tiling, and we use it to derive properties of the entire family. The CAP tiling has pure-point dynamical spectrum, which we compute explicitly, and comes from a natural cut-and-project scheme with $2$-dimensional Euclidean internal space. All other members of the Hat family, in particular the original version constructed from $30$-$60$-$90$ right triangles, are obtained via small modifications of the projection from this cut-and-project scheme.
Key words and phrases:
Tiling cohomology, Dynamical spectra, Model sets, Deformations, Monotile
2010 Mathematics Subject Classification:
52C20, 37D40, 55N05, 52C23
1. Introduction and results
In March 2023, David Smith, Joseph Myers, Craig Kaplan and Chaim Goodman-Strauss announced the construction of an aperiodic monotile called the Hat [17]. They showed that it is possible to tile the plane with isometric copies of this tile, but only non-periodically. In fact, the tilings utilize $12$ tiles up to translation: the original Hat, a reflected anti-Hat, and rotations of these tiles by multiples of $60$ degrees. There are essentially two kinds of tilings that result. In one, the Hats and anti-Hats assemble into four larger ‘meta-tiles’, called $\mathrm{T}$, $\mathrm{H}$, $\mathrm{F}$ and $\mathrm{P}$, which in turn form a fusion tiling. The meta-tiles assemble into larger ‘supertiles’ whose adjacencies are combinatorially the same as those of the basic meta-tiles. These assemble into second order supertiles, which assemble into third order supertiles, and so on to infinity. The other construction is similar, only using reflections of the $\mathrm{T}$, $\mathrm{H}$, $\mathrm{F}$ and $\mathrm{P}$ meta-tiles.
This result extended earlier results of Taylor [23] and of Socolar and Taylor [20]; see also [4, Ex. 6.6] for further details. There, a functional monotile is derived from an inflation rule with decorated hexagons that defines a unique local indistinguishability (LI) class of tilings with perfect local rules [4, Def. 5.20]. Unfortunately, these local rules are represented in the form of a decoration of a hexagon that encode both nearest-neighbor and next-to-nearest-neighbor information. A purely geometric version is possible, but only with a prototile that is not connected, hence not disk-like. Both the decorated hexagon and its reflected copy are needed; they occur with equal frequency, as is nicely visible from the Llama tiling; see [4, Fig. 6.23]. The unique LI class defined by the decorated hexagon is reflection symmetric, and the tiling is limit-periodic, with pure-point spectrum, both in the dynamical and in the diffraction sense.
The Hat achieves aperiodicity purely through geometry, with no need for any edge markings, using a tile that is a simple non-convex lattice polygon with $13$ edges. Once again, the reflected version of the Hat is also needed, but this time in such a way that reflection symmetry is broken. This leads to two distinct LI classes that are mirror images of one another (an enantiomorphic pair). That is, if we treat the Hat and the anti-Hat as two versions of the same tile, the local (matching) rules do not determine a single LI class, but a pair. It is possible to force a single LI class with local rules (for instance by declaring that two anti-Hats cannot touch), but then the rules for the Hat and anti-Hat are different and we have a tiling by two tile types. In this sense, the Hat construction satisfies the quest for a disk-like prototile that enforces aperiodicity, but at the price of slightly weakening the notion of local rules.
Figure 1. Patch of a Hat tiling. The anti-Hats are shaded. It has sixfold rotational symmetry, denoted by the cyclic group $C_{6}$, in the sense that each patch occurs in six orientations, but has no reflection symmetry.
An intriguing feature of the Hat construction is that it is possible to vary the relative lengths of the two kinds of edges in the Hat. The geometry continues to force the tiles to assemble into meta-tiles (or reflections of meta-tiles) with the same combinatorics as the original Hat tiling. After rescaling to preserve the average area per tile, and also rotating, the resultant tilings were observed to have approximately the same large-scale structure [17].
Our first main result is that the shape changes considered in [17] are part of a $4$-dimensional family of topological conjugacies between the translation dynamical systems of the family of Hat tilings. For the connections between them, we need the concepts of local derivability and mutual local derivability (MLD); see [4, Sec. 5.2] for definitions and background.
Theorem 1.
The moduli space of shape changes to the Hat tiling, modulo MLD equivalence, is $8$-dimensional. Four dimensions correspond to rigid linear transformations of $\mathbb{R}^{2}$. The other four induce topological conjugacies.
In particular, the translation dynamics of the $\mathbb{R}^{2}$ action on the hull of every version of a Hat tiling is topologically conjugate, up to linear transformation (possibly including reflection), to the translation dynamics on the hull of every other version of a Hat tiling. From a dynamical systems perspective, there is only one Hat tiling. Note that a Hat-dominated tiling is reflected into an anti-Hat dominated one, but since $\mathrm{GL}_{2}(\mathbb{R})$ is not connected, the two cannot continuously be deformed into one another.
Some of the shape changes violate the symmetries of the original construction. The family of shape changes that respect rotational symmetry is only $4$-dimensional and can be coded in two complex numbers, where we identify $\mathbb{R}^{2}$ with $\mathbb{C}$. One complex number parametrizes uniform rotations and rescalings, while the other parametrizes shape conjugacies. If we further insist that the shape changes respect reflectional symmetry, the two complex parameters must in fact be real. One represents rescalings and the other is exactly the ratio of edge lengths described in [17].
Within the family of shape deformations is a self-similar tiling, which we call the CAP tiling (because of its relation to the Hat and its simultaneous Cut-And-Project structure, proved below). With this set of shape parameters, the CAP tiling can be expressed as a geometric inflation tiling, not only a combinatorial one, although not one with a stone inflation. [^1] In a stone inflation, each super-meta-tile would have a footprint exactly $\phi^{2}$ larger than the original meta-tile, where $\phi=(1+\sqrt{5}\,)/2$ is the golden ratio (sometimes called $\tau$). However, the CAP tiling is MLD to a tiling by fractiles with an exact stone inflation. There are many tools available for studying self-similar tilings, especially those whose inflation factors are Pisot–Vijayaraghavan (PV) numbers, and we obtain the following result.
Theorem 2.
The dynamical spectrum of the CAP tiling is pure point. Up to scale and rotation, as detailed in Eq. (18) and Section 4, it is given by $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, where $\xi=\exp(\pi\hskip 0.5pt\mathrm{i}/3)$ is a primitive sixth root of unity. The CAP tiling itself is a cut-and-project tiling with internal space $\mathbb{R}^{2}$. With an appropriate choice of control points for the four meta-tiles, the total window is the hexagon shown in Figure 8, with subdivisions for the four types.
The projection structure of the CAP tiling, respectively its control point set, is robust under topological deformations as follows.
Theorem 3.
Every tiling that is topologically conjugate to the CAP tiling is MLD to a reprojection of the CAP tiling control points. That is, it is MLD to a cut-and-project set with the same total ((((or embedding))\,)) space and the same acceptance domain as the CAP tiling point set, only with a different projection from the total space to $\mathbb{R}^{2}$. Equivalently, it is MLD to a deformed model set with a linear deformation function on the window.
Here, topological conjugacy refers to the existence of a homeomorphism that commutes with the translation action of $\mathbb{R}^{2}$ on the tilings. Let us now turn to the topology of the continuous hull $\Omega_{\mathcal{T}}$ of a Hat tiling, where $\Omega_{\mathcal{T}}$ refers to the closure of the $\mathbb{R}^{2}$-orbit of the tiling in the standard local topology [4, 16]. The answer is the same for all choices of shape, and we use the standard Hat as a representative for all of them.
Theorem 4.
Let $\mathcal{T}$ be a Hat tiling and $\Omega_{\mathcal{T}}$ its continuous hull. Then, the Čech cohomology groups of $\Omega_{\mathcal{T}}$ are
[…]
This can further be decomposed into four real representations of the cyclic group $C_{6}$. This decomposition, and the way each component transforms under substitution, is given in Table 1.
Table 1. The integer cohomology of $\Omega_{\mathcal{T}}$ by representation and how it transforms under substitution.
Representation $r=1$ $r=\xi^{\pm 1}$ $r=\xi^{\pm 2}$ $r=-1$
$\check{H}^{0}(\Omega_{\mathcal{T}},\hskip 0.5pt\mathbb{Z})$hskip $\mathbb{Z}$ 0 0 0
Eigenvalues 1
$\check{H}^{1}(\Omega_{\mathcal{T}},\hskip 0.5pt\mathbb{Z})$ 0 $\hskip 0.5pt\mathbb{Z}^{4}$ 0 0
Eigenvalues $\phi^{\pm 2}$
$\check{H}^{2}(\Omega_{\mathcal{T}},\hskip 0.5pt\mathbb{Z})$ $\hskip 0.5pt\mathbb{Z}^{2}$ $\hskip 0.5pt\mathbb{Z}^{4}$ $\hskip 0.5pt\mathbb{Z}^{2}$ $\hskip 0.5pt\mathbb{Z}^{2}$
Eigenvalues $\phi^{\pm 4}$ $\phi^{\pm 2}$ $\xi^{\pm 1}$ $\phi^{\pm 2}$
The triangular lattice $\hskip 0.5pt\mathbb{Z}[\xi]$ appears in many places in this analysis. In particular, there are four special versions of the Hat, namely the Chevron, the Hat itself, the Turtle and the Comet, in which the vertices all lie on the lattice $c\hskip 0.5pt\hskip 0.5pt\mathbb{Z}[\xi]$ for some complex number $c$. The different values of $c$ are related by powers of $\sqrt{5}$, $\phi-\xi$ and $\phi-\xi^{5}$.
Remark 5.
The elements of $\hskip 0.5pt\mathbb{Z}[\xi]$ are the Eisenstein integers, which are the integers in the quadratic field $\mathbb{Q}(\sqrt{-3}\,)$, while the ring $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$ contains integers from the number field $\mathbb{Q}(\sqrt{-3},\sqrt{5}\,)$, but not all of them. This quartic field is Galois, with class number one, and has many interesting properties [^2] See entry $4.0.225.1$ of   https://www.lmfdb.org/NumberField   for some details., though we suppress a systematic use of them. $\Diamond$
The complex numbers $\phi-\xi^{5}$ and $\phi-\xi$ have norm $\sqrt{2}$ and argument ±$\pm\arctan\bigl{(}\sqrt{3/5\hskip 0.5pt}\,\bigr{)}$. Additional powers of $\phi-\xi^{5}$ and $\phi-\xi$ arise in the calculation of the dynamical spectrum. As a result, the spectra of the Chevron, Hat and Comet (but not the Turtle) are twisted relative to the lattices of these tilings. This twisting is different (in fact opposite) for the tiling by meta-tiles $\mathrm{T}$, $\mathrm{H}$, $\mathrm{P}$ and $\mathrm{F}$ and the tiling by the reflected meta-tiles. That is, the two different LI classes of the Chevron, Hat and Comet tilings have different spectra. A similar result holds for almost all choices of the Hat shape.
Theorem 6.
For all but two values of $\alpha/\beta$, where $\alpha$ and $\beta$ are the lengths of the two kinds of edges in the Hat, the spectra of the two LI classes of tilings built from Hats and anti-Hats in standard orientation are distinct.
When the spectrum of one LI class agrees with that of its reflected copy, we know that the corresponding dynamical systems are measure-theoretically isomorphic, by the Halmos–von Neumann theorem, which disregards sets of measure zero. However, this does not imply that the two systems are also topologically conjugate. In fact they are not for these two tilings, as we explain in Remark 15.
The organization of this paper is as follows. In Section 2, we consider shape deformations of the Hat. These are classified, up to MLD equivalence, by the vector-valued Čech cohomology group $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$, which we compute. We identify generators of $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$ with linear transformations and topological conjugacies, thereby proving Theorem 1. We then compute the integer-valued Čech cohomology of $\Omega_{\mathcal{T}}$, proving Theorem 4.
In Section 3, we construct the self-similar CAP tiling, whose topological dynamical system $(\Omega_{\mathcal{T}},\mathbb{R}^{2})$ is strictly ergodic by standard arguments. We show that it has pure-point dynamical spectrum, and we compute it explicitly. We then define control points and show how to view the resulting Delone set as a regular cut-and-project set, proving Theorem 2. Theorem 3 then follows from the cohomology calculations of Section 2.
In Section 4, we compare several versions of the Hat tiling, in particular the Chevron, the original Hat, the Turtle, the Comet and the CAP. These are all topologically conjugate, up to (complex) rescaling and rotation, where we compute the necessary rescalings, which all involve powers of $\phi-\xi$ and $\phi-\xi^{5}$. Combined with the results of Section 3, this yields the spectrum of the tilings obtained from each shape of monotile, proving Theorem 6. We close with some comments, observations and questions in Section 5.
2. Shape deformations of the Hat
Given any tiling in $d$ dimensions, we can study changes to the shapes and sizes of the tiles that allow the tiles to fit together in the exact same combinatorial patterns. Some of these changes result in tilings that are MLD to the original one. Infinitesimally, shape changes modulo MLD equivalence of a tiling $\mathcal{T}$ are parametrized by $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{d})$, where $\Omega_{\mathcal{T}}$ is the orbit closure (continuous hull) of the tiling $\mathcal{T}$; see [9]. Since $d=2$ and we identify $\mathbb{R}^{2}$ with $\mathbb{C}$, we must compute the complex-valued cohomology of the hull of each tiling. There are two such hulls, depending on whether the tiling involves meta-tiles or reflected meta-tiles. Since these hulls are homeomorphic (being reflections of one another), it suffices to work with a tiling by ordinary meta-tiles.
The proof of Theorem 1 proceeds by showing that $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{C})$ for the Hat tiling has complex dimension $4$ (hence real dimension $8$) and that the asymptotically negligible subspace corresponding to topological conjugacies has complex dimension $2$ (hence real dimension $4$). Most of this section is devoted to computing this cohomology.
One standard procedure for computing the cohomology of a substitution tilings was developed by Anderson and Putnam [1]. We construct a branched manifold $\varGamma$, called the Anderson–Putnam (AP) complex, from one representative of each type of tile in the tiling, with appropriate edge identifications. Specifically, whenever there is a place in the tiling where tiles of type $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ meet along an edge, the corresponding edges of $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ are identified in $\varGamma$. Substitution induces a map from this complex to itself. As long as the substitution meets a technical condition called forcing the border [11] (defined below), the tiling space is homeomorphic to the inverse limit of the AP complex under substitution, and the Čech cohomology $\check{H}^{k}(\Omega_{\mathcal{T}})$ of the tiling space is the direct limit of the (ordinary) cohomology $H^{k}(\varGamma)$ of the AP complex.
Anderson and Putnam developed their machinery for self-similar tilings, but the topology of the tiling space does not depend on the geometry of the tiles. A nearly identical procedure applies to combinatorial substitution systems like the Hat, or more generally to fusion tilings of any kind; see [10, Sec. 5] for details. The $n$th instance of the AP complex is replaced by a complex constructed from $n$th order supertiles identified along supertile edges and the substitution map (between a space and itself) is replaced by a forgetful map from a complex built from $n$th order supertiles to a complex build from $(n{\hskip 0.5pt-}1)$st order supertiles.
As long as the construction of $n$th order supertiles from $(n{\hskip 0.5pt-}1)$st order supertiles is combinatorially the same for every $n$, the complexes are all homeomorphic and have the same cohomology, and the forgetful maps induce the same maps on cohomology. Just as with the (naive) Anderson–Putnam procedure, the tiling space is the inverse limit of these complexes and the Čech cohomology $\check{H}^{k}(\Omega_{\mathcal{T}})$ of the tiling space is the direct limit of the cohomology $H^{k}(\varGamma)$ of the complexes.
A substitution (or fusion) tiling is said to force the border [11] if, for some $k$, every $k$th order supertile enforces a patch of tiles completely covering the supertile plus an extra margin of positive thickness around the supertile. If such a supertile is further inflated, this extra margin, where the tiling is also determined, grows in thickness at the same rate as the supertile. The result is that a single infinite-order supertile, if it does not already cover the whole plane, has a unique legal completion to the whole plane. It is this latter property we are interested in, as it simplifies the construction of the tiling space as an inverse limit space, and hence the computation of its Čech cohomology.
Figure 2. Patches of tiles enforced by the level-one supertiles. The tiles that belong to a supertile are shaded. These patches extend beyond the outline of the supertiles (thick lines) by a margin of positive thickness, thus proving the border-forcing property. There is some choice in the assignment of tiles to supertiles. Here, we use the original one from [17], while other choices are also possible, then leading to different fractalizations of tile edges (in comparison to Figure 3).
Lemma 7.
The substitution on the meta-tiles $\mathrm{T}$, $\mathrm{H}$, $\mathrm{P}$ and $\mathrm{F}$ forces the border.
Proof.
In Figure 2, we show patches of tiles which are enforced by the first order supertiles. The tiles intersecting the interior of their supertile were already given in [17, Figure 2.8]. We had to complement these patches only by a few tiles having a vertex on the supertile boundary. These extra vertices are either a purple tip of a tile or a supertile, where a local three-fold symmetry is required by the matching conditions. In the terminology of [4], we replace the inflation rule by a consistent pseudo-inflation with a complete ‘belt’ of positive thickness. These pictures show that the enforced patches extend by a positive margin beyond the perimeter of the supertile. Under inflation, the thickness of this margin, where the tiling is also determined, grows at the same rate as the supertile size, which implies border-forcing. ∎
Since the substitution forces the border, we can use the uncollared AP complex to compute the cohomology [1]. We work at the level of (meta-)tiles, but the construction for supertiles of any order is combinatorially identical.
Figure 3. First level supertiles with fractalized edges. These are obtained by iterating the inflation on a tile and rescaling the outline of the resulting patch back to the original size. For these fractiles, we have a stone inflation. The tiling obtained from this inflation is MLD with the CAP tiling from Figure 7.
Figure 4. The four meta-tiles in reference orientation. We have added markers to the edges as in [17] to fix the positive direction of each edge (except for the $L$ edge, which is non-directional, see below). In the positive direction, the marker is always on the left side of the edge.
In the Hat tiling, there are four kinds of meta-tiles, called $\mathrm{T}$, $\mathrm{H}$, $\mathrm{P}$ and $\mathrm{F}$, as shown in Figure 4. Each appears in six orientations, where we choose the reference orientation to be the one shown in Figure 4. That is, our tiles are $r^{k}\mathrm{T}$, $r^{k}\mathrm{H}$, $r^{k}\mathrm{P}$ and $r^{k}\mathrm{F}$, where $k$ ranges from 00 to $5$ and $r$ means ‘positive rotation by $60$ degrees’. The inflations of these tiles can be read off from Figure 2.
Figure 5. The five types of (oriented) edges $A$, $B$, $X$, $F$ and $L$ in reference orientation, along with their inflations. In the positive edge direction, the marker is on the left (except for the $L$ edge, which is non-directional). The edge inflations can in principle be extracted from Figure 2, but the determination of edge orientations is a bit subtle.
There are five kinds of edges, labeled $A$, $B$, $X$, $F$, and $L$, all appearing in Figure 5. The first four edges have a clear direction. We define the reference orientation to be vertical (or roughly vertical for $F$), with the relevant marking on the left. The $L$ edge is not directional, so as a chain we have $r^{3}L=-L$.
Now, we turn to vertices. Let $A_{+}$ and $A_{-}$ denote the endpoints and beginning points of an $A$ edge, and likewise with $B$, $X$, $F$ and $L$. The symmetry of the $L$ edge implies that $L_{+}=r^{3}L_{-}$. Other relations are obtained from Figure 4 by identifying the end of one edge with the beginning of the next edge on the boundary of a meta-tile. This yields the following relations:
(1) $A_{+}$ $\,=\,r^{5}X_{-},$ $A_{-}$ $\,=\,r^{3}X_{-},$
(2) $B_{+}$ $\,=\,X_{-},$ $B_{-}$ $\,=\,r^{2}X_{-},$
(3) $X_{+}$ $\,=\,X_{+},$ $X_{-}$ $\,=\,X_{-},$
(4) $F_{+}$ $\,=\,X_{+},$ $F_{-}$ $\,=\,F_{-},$
(5) $L_{+}$ $\,=\,X_{-},$ $L_{-}$ $\,=\,r^{3}X_{-},$
expressing every vertex in terms of $X_{-}$, $F_{-}$ and $X_{+}$. Furthermore,
(6) $F_{-}\,=\,r^{2}F_{-}\,=\,r^{4}F_{-}\hskip 0.5pt,\qquad\qquad X_{+}\,=\,r^{2}X_{+}\,=\,r^{4}X_{+}\hskip 0.5pt.$
Having identified the faces, edges and vertices of the AP complex, we now compute the boundary maps, relative to the bases $\{\mathrm{T},\mathrm{H},\mathrm{P},\mathrm{F}\}$ for faces, $\{A,B,X,F,L\}$ for edges and $\{X_{-},X_{+},F_{-}\}$ for vertices.
(7) $\partial_{1}\,=\,\begin{pmatrix}r^{5}{-}\hskip 0.5ptr^{3}&1{-}\hskip 0.5ptr^{2}&-1&0&1{-}\hskip 0.5ptr^{3}\\ 0&0&1&1&0\\ 0&0&0&-1&0\end{pmatrix}$
(8) $\partial_{2}\,=\,\begin{pmatrix}r{+}\hskip 0.5ptr^{5}&-1&r^{5}&0\\ -1&r{+}\hskip 0.5ptr^{5}&-r^{5}&-r^{5}\\ 0&0&(r{-}\hskip 0.5ptr^{5})(1{+}\hskip 0.5ptr^{3})&r^{4}{-}\hskip 0.5ptr^{2}\\ 0&0&0&r\hskip 0.5pt{-}\hskip 0.5ptr^{3}\\ 0&0&0&r^{4}{+}\hskip 0.5ptr^{5}\end{pmatrix}.$
Here, we view chains as column vectors, to be acted on from the left by $\partial_{1}$ and $\partial_{2}$. Co-chains are row vectors, to be acted on from the right. Strictly speaking, each entry should be viewed as a $6\times 6$ block, with ‘1’ being the identity matrix and
$r\,=\,\begin{pmatrix}0&0&0&0&0&1\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\end{pmatrix},$
and we further need to apply the identities $L=-r^{3}L$, $X_{+}=r^{2}X_{+}$ and $F_{-}=r^{2}F_{-}$. Fortunately, it is not necessary to work with gigantic matrices. Instead, we follow the methods of [15], see also [16, Ch. 4], and work with one representation of $C_{6}$ at a time. The total cohomology is then the direct sum of the contributions of each representation.
Working over the complex numbers, a representation is just a choice of a $6$th root of unity to assign to $r$. That is, we must consider $r=\xi^{k}$ where $k\in\{0,1,2,3,4,5\}$. However, the rows corresponding to $X_{+}$ and $F_{-}$ only appear in the representations where $r^{2}=1$, and the row and column corresponding to $L$ only appear in the representations where $r^{3}=-1$.
Remark 8.
Computing the real-valued and integer-valued cohomology is somewhat more complicated, both because the irreducible real representations of $C_{6}$ are not all $1$-dimensional and because, for integer-valued cohomology, decomposing into representations is only guaranteed to compute a finite-index subgroup of the cohomology. We will compute the integer-valued cohomology of the Hat tiling later in this section. $\Diamond$
The sizes and ranks of the matrices $\partial_{1}$ and $\partial_{2}$ in each representation are shown in Table 2, together with the contributions of each representation to $H^{0}(\varGamma,\mathbb{C})$, $H^{1}(\varGamma,\mathbb{C})$ and $H^{2}(\varGamma,\mathbb{C})$.
Table 2. The contribution of each representation of $C_{6}$ to the complex cohomology $H^{k}(\varGamma,\mathbb{C})$ of the Anderson–Putnam complex
Representation $r=1$ $r=\xi$ $r=\xi^{2}$ $r=\xi^{3}$ $r=\xi^{4}$ $r=\xi^{5}$
Number of faces 4 4 4 4 4 4
Number of edges 4 5 4 5 4 5
Number of vertices 3 1 1 3 1 1
Rank of $\partial_{2}$ 2 2 3 2 3 2
Rank of $\partial_{1}$ 2 1 1 3 1 1
[…]
Contribution to $H^{1}$ 0 $\mathbb{C}^{2}$ 0 0 0 $\mathbb{C}^{2}$
[…]
Having computed $H^{1}(\varGamma,\mathbb{C})$, we must take a direct limit under the pullback of the forgetful map to obtain the Čech cohomology $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{C})$ of the hull $\Omega_{\mathcal{T}}$. There are manifestly two generators of $H^{1}(\varGamma,\mathbb{C})$ that transform with eigenvalue $\phi^{2}$ under substitution. This is because the cochains that assign to each edge its horizontal or vertical displacement get stretched asymptotically by $\phi^{2}$ under substitution. Since substitution acts on $1$-cochains, and therefore on $H^{1}(\varGamma,\mathbb{C})$, via a matrix with integer coefficients, the algebraic conjugate of $\phi^{2}$, namely $\phi^{-2}$, must also be an eigenvalue with multiplicity two. Since $H^{1}(\varGamma,\mathbb{C})$ is only $4$-dimensional, substitution acts invertibly on $H^{1}(\varGamma,\mathbb{C})$ and the direct limit of $H^{1}(\varGamma,\mathbb{C})=\mathbb{C}^{4}$ is just $\mathbb{C}^{4}$. We conclude that $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{C})=\mathbb{C}^{4}=\mathbb{R}^{8}$, with a subspace of real dimension $4$ that expands under substitution with eigenvalue $\phi^{2}$ and a subspace of real dimension $4$ that contracts under substitution with eigenvalue $\phi^{-2}$.
The set of asymptotically negligible classes, which describe shape changes (modulo MLD) that are topological conjugacies, is exactly the contracting subspace under substitution [9]. Thus, the $8$-dimensional family of shape deformations of the Hat (modulo MLD) breaks into a $4$-dimensional family of changes to the large-scale structure and a $4$-dimensional family of shape conjugacies.
Since we can transform any tiling by applying a rigid linear transformation, thereby changing its large-scale structure, and since the space of linear transformations is $4$-dimensional, the shape classes of any two Hat tilings have expansive components that are related by such a linear transformation. After applying a linear transformation to one of them, the shape classes differ by something asymptotically negligible, so the tiling spaces are topologically conjugate.
So far, the analysis has only applied to tilings by meta-tiles, not by anti-meta-tiles. However, tilings by anti-meta tiles are already related by a linear transformation (namely reflection) to tilings by standard meta-tiles. Thus, every space of tilings by Hats, whether involving meta-tiles or anti-meta-tiles, is topologically conjugate, up to linear transformation, to every other such space. [^3] Note that a Hat tiling is homeomorphic, but not topologically conjugate, to an anti-Hat tiling. Likewise, this applies also to the shape changes that are used to define Socolar’s Key tiles [19].
This concludes the proof of Theorem 1.
Remark 9.
Both the expansive classes and the asymptotically negligible classes come equally from the $r=\xi$ and $r=\bar{\xi}=\xi^{5}$ representations. If we want to preserve rotational symmetry, the vector along the $r^{k}E$ edge (where $E$ is any of $\{A,B,X,F,L\}$) should be $\xi^{k}$ times the vector along the $E$ edge. That is, shapes that respect rotational symmetry live entirely in the $r=\xi$ representation, while contributions from the $r=\xi^{5}$ representation break rotational symmetry. $\Diamond$
A (relatively) simple computation of $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{C})$ was sufficient to prove Theorem 1. However, a more complete computation of the cohomology of $\Omega_{\mathcal{T}}$, and how it transforms under rotation and substitution, is also of interest, yielding insight into the structure of Hat tilings.
First, we consider real-valued cohomology. For any topological space $X$, any cohomology theory, and any degree $k$, the real dimension of $H^{k}(X,\mathbb{R})$ is the same as the complex dimension of $H^{k}(X,\mathbb{C})$. This implies that $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R})=H^{1}(\varGamma,\mathbb{R})=\mathbb{R}^{4}$ and that $\check{H}^{2}(\Omega_{\mathcal{T}},\mathbb{R})=H^{2}(\varGamma,\mathbb{R})=\mathbb{R}^{10}$. The decomposition into representations is slightly more subtle, since the complex representations with $r$ being $\xi$, $\xi^{2}$, $\xi^{4}$ or $\xi^{5}$ are not complexifications of real representations. Rather, the direct sum of the $r=\xi^{\pm 1}$ representations is the complexification of a $2$-dimensional real representation, as is the direct sum of the $r=\xi^{\pm 2}$ representations.
Rather than simply relying on stretching and algebraic conjugacy, we compute directly how substitution acts on $H^{1}(\varGamma,\mathbb{R})$ and $H^{2}(\varGamma,\mathbb{R})$. The action of substitution on faces can be read off from Figure 6 and the action on edges and vertices from Figure 5. The matrices for the action on vertices, edges, and faces are:
[…]
Acting on $H^{2}(\varGamma,\mathbb{R})$, we can study this one representation at a time. The substitution eigenvalues are exactly those listed in Table 1. Note that the product of these eigenvalues is 1, so substitution acts via a unimodular matrix on $H^{2}(\varGamma,\mathbb{R})$.
We next compute $H^{*}(\Omega_{\mathcal{T}},\hskip 0.5pt\mathbb{Z})$ more carefully, without decomposing by representation, using the full $27\times 24$ matrix $\partial_{2}$ and the full $10\times 27$ matrix $\partial_{1}$. The result is that $H^{1}(\varGamma,\hskip 0.5pt\mathbb{Z})=\hskip 0.5pt\mathbb{Z}^{4}$ and $H^{2}(\varGamma,\hskip 0.5pt\mathbb{Z})=\hskip 0.5pt\mathbb{Z}^{10}$. Since substitution acts via unimodular matrices, the direct limits of $\hskip 0.5pt\mathbb{Z}^{4}$ and $\hskip 0.5pt\mathbb{Z}^{10}$ are simply $\hskip 0.5pt\mathbb{Z}^{4}$ and $\hskip 0.5pt\mathbb{Z}^{10}$. As with the real-valued cohomology, these can be decomposed into direct sums of modules of the form $\hskip 0.5pt\mathbb{Z}[r]/(r-1)$ (corresponding to $r=1$), $\hskip 0.5pt\mathbb{Z}[r]/(r+1)$ (corresponding to $r=-1$), $\hskip 0.5pt\mathbb{Z}[r]/(r^{2}-r+1)$ (corresponding to $r=\xi^{\pm 1}$) and $\hskip 0.5pt\mathbb{Z}[r]/(r^{2}+r+1)$ (corresponding to $r=\xi^{\pm 2}$). This establishes Theorem 4.
Finally, we turn to a closer examination of $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{C})\simeq H^{1}(\varGamma,\mathbb{C})$, which has components in the $r=\xi^{\pm 1}$ representations. In the $r=\xi$ representation, the left-kernel of the matrix $\partial_{2}$ is spanned by the three row vectors
(12) $\begin{split}\boldsymbol{\mu}_{1}&\,=\,(1+\xi^{5},1+\xi^{5},1,0,0)\hskip 0.5pt,\\ \boldsymbol{\mu}_{2}&\,=\,(\xi+\xi^{2},\xi+\xi^{2},0,1,0)\hskip 0.5pt,\\ \boldsymbol{\mu}_{3}&\,=\,(1+\xi^{5},1+\xi^{5},0,0,1)\hskip 0.5pt.\end{split}$
The row space of $\partial_{1}$ is spanned by $2\boldsymbol{\mu}_{3}-\boldsymbol{\mu}_{1}$, so $\boldsymbol{\mu}_{1}$ is cohomologous to $2\boldsymbol{\mu}_{3}$.
Under right-multiplication by $\sigma_{1}$, $\boldsymbol{\mu}_{1}$ transforms to $(2+\xi^{5})\boldsymbol{\mu}_{1}+2\boldsymbol{\mu}_{2}$, while $\boldsymbol{\mu}_{2}$ transforms to $\xi(\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2})$. That is, the action of substitution on $H^{1}(\varGamma,\mathbb{C})$ in the $\{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}\}$ basis and in the $r=\xi$ representation is via the matrix
(13) $\widetilde{\sigma}_{1}\,=\,\begin{pmatrix}2{+}\hskip 0.5pt\xi^{5}&2\\ \xi&\xi\end{pmatrix},$
where as usual we are thinking of cohomology classes as rows with the matrix acting to the right. The left and right eigenvectors corresponding to the eigenvalues $\lambda=\phi^{\pm 2}$ take the form
(14) $\ell_{\lambda}\,=\,(\lambda-\xi,2),\qquad r_{\lambda}\,=\,\begin{pmatrix}\lambda-\xi\\ \xi\end{pmatrix}.$
Results for $r=\xi^{5}$ are complex conjugates of the corresponding results for $r=\xi$.
Figure 6. The tiles of the CAP tiling, together with their control points (orange, blue, green and purple for $\mathrm{T}$-, $\mathrm{H}$-, $\mathrm{P}$- and $\mathrm{F}$-tiles, respectively). Relative to the corners marked with black dots, the control points are at positions $\hskip 0.5pt\mathrm{i}\phi\xi$, $-\hskip 0.5pt\mathrm{i}\phi\xi$, $\hskip 0.5pt\mathrm{i}\phi$ and $\hskip 0.5pt\mathrm{i}\phi$, respectively. Apart from the $\mathrm{H}$-tiles, the control points are outside their tile, but never coincide with a control point of a neighboring tile. Each tile occurs in six different orientations, but not in a reflected version.
3. The CAP tiling
In this section, we construct the CAP tiling. We work directly with the shape of the $\mathrm{T}$, $\mathrm{H}$, $\mathrm{P}$ and $\mathrm{F}$ meta-tiles, or equivalently the vector displacements along the $A$, $B$, $X$, $F$ and $L$ edges. We choose
$(A,B,X,F,L)\,=\,(3\phi,3\phi,1,\phi+\xi,2\phi)\,=\,\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}+2\phi\boldsymbol{\mu}_{3}\hskip 0.5pt.$
This is cohomologous to $(\phi+1)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}$, which we represent as $(\phi+1,\phi+\xi)$. This is a left-eigenvector of $\widetilde{\sigma}_{1}$ with eigenvalue $\phi^{2}$. The substitution rule (shown in Figure 2) is self-similar in the sense that a substituted tiling (with shape class $(\phi+1,\phi+\xi)\widetilde{\sigma}_{1}$) is then MLD to a rescaling of the original tiling by $\phi^{2}$ (with shape class $\phi^{2}(\phi+1,\phi+\xi)$). This constitutes the LIDS property in the sense of [4, Def. 5.16].
Figure 7. A patch of the CAP tiling. All control points occur in triples inside the $\mathrm{H}$-tiles. One point (blue) of such a triple comes from the $\mathrm{H}$-tile itself, the other two from the two tiles adjacent along one of the two blue edges. The distances between any control points are elements of the return module $R$.
The tiles with these shape parameters are shown in Figure 6, where we have also added a control point for each tile. A patch of the resulting tiling, also including control points, is shown in Figure 7. The control points have been chosen such that they lie in a single orbit of the return module $R$ of the tiling. The return module is the $\hskip 0.5pt\mathbb{Z}$-span of all return vectors, which in turn are vectors which translate a finite patch of a tiling to an identical copy of that patch in the same tiling.
Control points of tiles of the same type and orientation are clearly in a single $R$-orbit, but if this is to hold for all control points, a number of constraints must be satisfied. In Figure 7, one can check that
[…]
is a return vector, and that all other return vectors are contained in the module generated by $\phi\hskip 0.5ptt_{0}$, $t_{0}$, $\phi\hskip 0.5pt\xi t_{0}$, and $\xi t_{0}$. This module is $t_{0}\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, which is invariant under inflation, and is indeed our return module $R$ (see also Section 4). In Figure 7, one can also convince oneself that all control points are separated by vectors in $R$.
The return module $R$ is needed to determine the point spectrum of the CAP tiling, but also to represent its set of control points as a cut-and-project set. The latter is possible only if the dynamical spectrum is pure point, which we show next. A classical method to prove this for self-similar inflation tilings is Solomyak’s overlap algorithm [21], which works as follows. Suppose $\mathcal{T}$ is an instance of such an inflation tiling. Solomyak showed that its spectrum is pure point if and only if, for every return vector $t$,
[…]
where $\sigma$ is the inflation map, and $\mathrm{coinc}(\mathcal{T}_{1},\mathcal{T}_{2})$ is the area fraction covered by coincident tiles of the two tilings $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. It is actually enough to show this for $t$ in a basis of the return module. The area not covered by coincident tiles is called the discrepancy area.
The space can now be cut into compact pieces, called overlaps, which are just the (non-empty) overlaps of all pairs of tiles, one from each tiling. Overlaps formed by two coincident tiles (coincidence overlaps) cover the coincidence area, whereas the remaining discrepancy overlaps cover the discrepancy area. There is an inflation defined on overlaps: one inflates the two tiles, and decomposes the overlap of the two supertiles into overlaps. Due to finite local complexity, for any fixed return vector $t$, the number of different overlap types (up to translation) is finite, and a PV inflation factor guarantees that only a finite number of overlap types is generated under inflation.
For a stone inflation, the corresponding overlap inflation can now be used to test Solomyak’s criterion. The area fraction covered by coincident tiles converges to $1$ if and only if every overlap type eventually produces a coincidence overlap. In that case, the area fraction covered by the discrepancy overlaps shrinks at each inflation step by a certain factor bounded away from $1$. Conversely, if at least one overlap type does not produce any coincidence, there is a discrepancy region whose area fraction does not shrink.
Here, we have a stone inflation only for the fractiles, for which it is difficult to determine whether two tiles have a non-empty overlap. For this situation, there is a refined algorithm by Akiyama and Lee [2]. It works with potential overlaps in a first stage, and eliminates those which are not true overlaps at a later stage. This refined algorithm shows that the CAP tiling has pure point spectrum. Unfortunately, the details are complicated and had to be done by computer. For readers who find computer-assisted proofs unsatisfying, we give an alternative proof of pure-point spectrum below.
Lemma 10.
The topological dynamical system $(\Omega_{\mathcal{T}},\mathbb{R}^{2})$ induced by the CAP tiling is strictly ergodic. Further, it has pure-point dynamical spectrum, where all eigenfunctions have continuous representatives.
Figure 8. Window for the cut-and-project scheme of the CAP tiling. The color code refers to the type of the tile whose control point is plotted (orange for $\mathrm{T}$-, blue for $\mathrm{H}$-, green for $\mathrm{P}$- and purple for $\mathrm{F}$-tiles). The subdivision for the different tile types is partly fractal, see Eq. (17), and partly straight. The edge length of the window is $\phi$ times the unit length in $R$. The image is chiral with sixfold rotational, but no reflection symmetry, although the total window, and hence the set of all control points (without color), is mirror symmetric.
Now, we show that the set $\varLambda$ of control points of the CAP tiling, which is a subset of the return module $R$, forms a cut-and-project set (CPS). For this, we need the construction from [8] which gives the CPS from intrinsic data of the tiling; see also [22, Sec. 5.2] for a detailed description. The procedure simplifies for inflation point sets, where one can more easily determine the limit translation module from the return vectors. The upshot is that the limit translation module agrees with our return module, and needs to be lifted to an embedding lattice. Due to the underlying structure with the PV unit $\phi^{2}$, one natural way to turn $R$ into a lattice in $\mathbb{R}^{4}$ employs the classic Minkowski embedding; see [4, Sec. 3.4].
To this end, we first rescale $\varLambda$ by a factor $t^{-1}_{0}$, see Eq. (15), so that its return module is equal to $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$. This subset is then easily lifted to the Minkowski embedding of $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, and then projected to internal space. The Minkowski embedding $\mathcal{L}$ of $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$ is obtained by combining the basis vectors $\{\phi,1,\xi\phi,\xi\}\subset\mathbb{C}$ with their Galois conjugates, so that we obtain the basis vectors of $\mathcal{L}$ as $\{(\phi,1-\phi),(1,1),(\phi\hskip 0.5pt\xi,(1-\phi)\xi^{5}),(\xi,\xi^{5})\}\subset\mathbb{C}^{2}$. The resulting window is shown in Figure 8. This is justified by the general uniform distribution result for this setting [14].
Remark 11.
The choice of control points is not unique, and any other representative from the same MLD class can be used as well. Matters are simplified by using only one translation class with respect to $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, as one has to use several windows for the different translation classes otherwise (as also occurs in [19]). Even so, other choices are possible with simple total window, for instance one where the center of the outer hexagon is a smaller hexagon for all control points of $\mathrm{F}$-tiles, surrounded by a belt of six trapezoids that are fractally subdivided into the other three types, which then breaks the reflection symmetry. This choice would simply amount to a re-coloring of the control points we use here. $\Diamond$
Put differently, we are using the following cut-and-project scheme [4, Sec. 7.2]
[…]
with the natural projections $\pi$ and $\pi_{\mathrm{int}}$ to direct (or physical) and internal (or perpendicular) space. The $\star$-map is the Galois automorphism induced by the joint action of $\phi\mapsto\phi^{\hskip 0.5pt\prime}=1-\phi$ and $\xi\mapsto\bar{\xi}=\xi^{5}$. This is the mapping used for the Minkowski embedding. It is not the only choice, but any other one is equivalent to it via standard modifications.
The point density of our CPS with the above lattice and window is given by the density of the lattice times the area of the total window. The lattice $\mathcal{L}$ has a unit cell of volume $15/4$, so we get $\mathrm{dens}(\mathcal{L})=4/15$. Indeed, each summand $\hskip 0.5pt\mathbb{Z}[\phi]$ provides a factor $\sqrt{5}$ (as known from the Fibonacci tiling), and because the sum is not orthogonal, we get two extra factors $\sqrt{3}/2$, as known from the triangular lattice. The window is a regular hexagon of edge length $\phi$, hence has area $\frac{3}{2}\phi^{2}\sqrt{3}$, so that the density of the covering model set with this window becomes
$\rho_{1}\,=\,\frac{\raisebox{-2.0pt}{$2$}}{\raisebox{0.5pt}{$5$}}\sqrt{3}\,\phi^{2}.$
Let us now describe in more detail the cut-and-project setting in terms of the inflation. For this purpose, denote by $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{24}$ the $24$ translational prototiles, which emerge from the $6$ orientations of the tiles $\mathrm{T}$, $\mathrm{H}$, $\mathrm{P}$ and $\mathrm{F}$. Let $D=(D_{ij})$ be the displacement matrix, where $D_{ij}$ is the set of relative positions of tiles of type $j$ in supertiles of type $i$, as determined via the control points. With $\varLambda_{i}$ denoting the control points of type $i$ in a fixed point tiling (for instance the one generated by a hexagon with its control point at the origin), one then has
$\varLambda_{i}\,=\>\bigcup_{j}^{.}\phi^{2}\varLambda_{j}\hskip 0.5pt+D_{ij}\hskip 0.5pt,$
where $A+B$ denotes the Minkowski sum of two sets, and the unions on the right-hand side are disjoint.
Under the $\star$-map, and upon taking closures, one obtains the window iterated function system (IFS)
$W_{i}\,=\>\bigcup_{j}\phi^{-2}W_{j}\hskip 0.5pt+D_{ij}^{\star}$
for $W_{i}=\overline{\varLambda_{i}^{\star}}$. Note that, due to having taken the closure, the unions on the right-hand side need no longer be disjoint. Since $\phi^{2}$ is a PV number, this IFS is contractive and defines a unique attractor $(W_{1},\ldots,W_{24})\subset(\mathcal{K}\mathbb{R}^{2})^{24}$, where $\mathcal{K}\mathbb{R}^{2}$ is the space of non-empty compact subsets of $\mathbb{R}^{2}$, equipped with the Hausdorff metric. By construction, each
$(W_{i})\,\mathrel{\mathop{:}}=\,\{x\in\hskip 0.5pt\mathbb{Z}[\xi,\phi]:x^{\star}\in W_{i}\}$
is a regular model set with $\varLambda_{i}\subseteq\mbox{\Large$\curlywedge$}(W_{i})$ by construction, where each window $W_{i}$ is a topologically regular compact set with almost no boundary; see [6, Sec. 3] and references therein for details. Some of the boundaries are straight lines, while some others are fractals of Koch curve type, then with Hausdorff dimension
[…]
as determined by standard methods, see Figure 9, from the induced edge IFS.
Figure 9. Koch curve type construction of the fractal part of the window boundaries, for the $\mathrm{H}$-tile (left) and the $\mathrm{P}$-tile (right). Here, every occurring corner is a projected lattice point, so the fractal curves contain lots of them.
We now provide the alternate proof that the CAP tiling has pure-point dynamical spectrum. We will show that the $\varLambda_{i}$ have the same density as the model sets $(W_{i})$, thereby proving pure-point diffraction spectrum and hence pure-point dynamical spectrum via the equivalence theorem from [12, 7]. Since the $\varLambda_{i}$ are disjoint, it suffices to show that
$\mathrm{dens}(\varLambda_{1}\cup\ldots\cup\varLambda_{24})\,=\,\mathrm{dens}\bigl{(}\mbox{\Large$\curlywedge$}(W)\bigr{)}$
for $W=W_{1}\cup\ldots\cup W_{24}$, the window of Figure 8, where we know that the $W_{i}$ are mutually interior-disjoint.
The value of the density of the set of tiles or control points of the inflation tiling can also be computed as follows. The relative frequencies of the four types of tiles are obtained from the right Perron–Frobenius (PF) eigenvector of the meta-tile inflation matrix. Normalized such that these frequencies add up to $1$, we get for the tiles $(\mathrm{T},\mathrm{H},\mathrm{P},\mathrm{F})$ the frequency vector
$\boldsymbol{f}\,=\,\bigl{(}\tfrac{5}{3}-\phi,\tfrac{1}{3},2\,\phi-3,2-\phi\bigr{)},$
which is independent of the choice of inflation. We remark that one third of all tiles are hexagons. For a stone inflation, the relative areas of the four tile types could be extracted from the entries of the left PF eigenvector. Here, however, these would be the relative areas of the corresponding fractiles, whose absolute areas are not immediately known. We therefore take the areas of the polygonal tiles instead. With $a=\sqrt{3}/4$, the area of an equilateral triangle of edge length $1$, we obtain, after some algebra, the vector of tile areas
$\boldsymbol{v}\,=\,a\bigl{(}9+9\,\phi,15+27\,\phi,14+22\,\phi,15+23\,\phi\bigr{)}.$
For this computation, we have used the lengths of the tile edges according to
$(A,B,X,F,L)\,=\,(3\hskip 0.5pt\phi,3\hskip 0.5pt\phi,1,\phi+\xi,2\hskip 0.5pt\phi)\hskip 0.5pt.$
The average area per tile then becomes $\boldsymbol{v}\cdot\boldsymbol{f}$, and its inverse is the tile density. However, we first want to rescale the tiling such that the generator $3\phi+2-\xi$ of the return module $R$ has unit length, so that we have to multiply the density by $|3\phi+2-\xi|^{2}=6(2+3\phi)$. At this new scale, we obtain a tile density
$\rho_{2}\,=\,\frac{|3\phi+2-\xi|^{2}}{\boldsymbol{v}\cdot\boldsymbol{f}}\,=\,\frac{\raisebox{-2.0pt}{$2$}}{\raisebox{0.5pt}{$5$}}\sqrt{3}\,\phi^{2}.$
Since $\rho_{1}=\rho_{2}$, the set of control points of the inflation tiling is indeed the model set $(W)$, possibly up to a difference set of zero density. The autocorrelation measures of both sets agree, because the addition or subtraction of sets of zero density does not matter [4, Rem. 9.14], and the same applies to the diffraction of Dirac combs on $(W)$ that are weighted according to the point classes, in comparison to the correspondingly weighted Dirac combs on the control point sets. The diffraction measures of both situations are thus the same, and pure point. Since the hulls of both sets give rise to ergodic dynamical systems, the equivalence theorem [12, 7] implies that the tiling dynamical system has pure-point dynamical spectrum, and both systems have eigenfunctions with continuous representatives. In other words, topological and measure-theoretic point spectra are the same.
Remark 12.
The equality of $\rho_{1}$ and $\rho_{2}$ is also a consistency check of the correct choice of lattice for the CPS. The result now gives the group $\mathbb{R}^{4}/\mathcal{L}\simeq\mathbb{T}^{4}$ as the maximal equicontinuous factor (MEF) of our tiling space. On the level of the control points, due to the nature of the windows, the sets $\varLambda_{i}$ and $(W_{i})$ can only differ in points whose $\star$-image lies on a window boundary. If this happens (as it does here), the model set description gives the union of all elements in the fibre over the corresponding point of the MEF. They are the singular points, which have a slightly more complicated nature than usual, due to the partly fractal nature of the window boundaries. $\Diamond$
The dynamical spectrum now is the projection of $\mathcal{L}^{*}$, the dual lattice of the Minkowski embedding $\mathcal{L}$ of $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, into direct space, which is the dual module. Here, one finds
[…]
by a standard calculation, which gives a scaled and rotated copy of $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$ as claimed.
This concludes the proof of Theorem 2 and also that of Lemma 10.
Remark 13.
The total window for all control points is a regular hexagon, whereas the window for the $\mathrm{F}$-tile control points consists of $6$ regular triangles. By the general criteria for MLD relations of regular model sets in the same CPS, see [4, Rem. 7.6], we thus see that the $\mathrm{F}$-type points can locally be derived from knowing all points (without color), and vice versa. All other subwindows partly have fractal boundaries. This means that control points of other tile types cannot be derived by local means from the (uncolored) positions of all control points. $\Diamond$
Let us say a bit more on the diffraction measure, formulated for the control point set of the CAP tiling. Each of the $24$ Delone sets $\varLambda_{i}$ gives rise to a Dirac comb $\delta_{\!\varLambda_{i}}\mathrel{\mathop{:}}=\sum_{x\in\varLambda_{i}}\delta_{x}$, and thus to the corresponding Fourier–Bohr coefficient
$a_{i}(k)\,\mathrel{\mathop{:}}=\lim_{s\to\infty}\frac{\raisebox{-2.0pt}{$1$}}{\raisebox{0.5pt}{$\mathrm{vol}(B_{s}(0))$}}\sum_{x\in\varLambda_{i}\cap B_{s}(0)}\hskip 0.5pt\mathrm{e}^{-2\pi\hskip 0.5pt\mathrm{i}kx},$
which exists, for any $k\in\mathbb{R}^{2}$, due to ergodicity. Here, $B_{s}(0)$ denotes the closed disk of radius $s$ around 00. The Fourier–Bohr coefficient is an area-averaged exponential sum, with $a_{i}(0)=\mathrm{dens}(\varLambda_{i})=\mathrm{dens}(\mathcal{L})\hskip 0.5pt\mathrm{vol}(W_{i})$ from our above analysis. By the general spectral result for regular model sets, compare [4, Thm. 9.4], we get
(19) $a_{i}(k)\,=\,\frac{\mathrm{dens}(\varLambda_{i})}{\mathrm{vol}(W_{i})}\,\widehat{1_{W_{i}}}(-k^{\star})$
for any $k$ in the dynamical spectrum, and $a_{i}(k)=0$ otherwise. Here, $1_{A}$ denotes the characteristic function of the set $A$, and $\widehat{1_{A}}$ its Fourier transform.
If we now consider the weighted Dirac comb $\omega=\sum_{i}h_{i}\hskip 0.5pt\delta_{\!\varLambda_{i}}$, which is pure-point diffractive, one obtains the diffraction measure
$\widehat{\gamma}\>=\sum_{k\in L^{\circledast}}\big{\lvert}h_{1}a_{1}(k)+\ldots+h_{24}a_{24}(k)\big{\rvert}^{2}\delta_{k}\hskip 0.5pt,$
which is a pure-point measure supported on the Fourier module $L^{\circledast}=\pi(\mathcal{L}^{*})$, with $\pi$ as in Eq. (16). The module $L^{\circledast}$ equals the dynamical spectrum from above. The validity of this formula follows from the fact that a primitive inflation tiling leads to Delone sets that have the phase consistency property [4, 3, 6]. With little effort, this can be generalized to other choices of control points, and to convolutions with various profile functions.
Remark 14.
While the formula in (19) is a major result in the theory of regular model sets, its calculation can become complicated, in particular for windows with fractal boundaries. Here, the windows $W_{i}$ are still such that a numerical approach (as in [19]) gives reasonable approximations. An exact calculation is also possible, via the compactly converging Fourier cocycle developed in [6, Sec. 4], which works here in exactly the same way. A simple example, also in comparison to a numerical approximation, is worked out in [5]. $\Diamond$
This is the standard approach for regular model sets, which directly applies to the CAP tiling and its control points. To transfer any of this to the other members of the Hat family of tilings, we need to connect them to the same cut-and-project scheme, as we do next.
3.1. Reprojections
Imagine a Delone set produced like the control points of the CAP tiling, with exactly the same total space and exactly the same window, only with a different projection from the total space to physical space. For this, we need some notation. If $z\in\mathcal{L}$ is an element of our embedding lattice, we write it as $z=(x,x^{\star})$, where $x=\pi(z)$ and $x^{\star}=\pi_{\text{int}}(z)$ are the orthogonal projections of $z$ into direct (or physical) and internal space, respectively.
If $L$ is an arbitrary linear map from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$, then
(20) $\pi^{\prime}(z)\,=\,\bigl{(}\pi+L\circ\pi_{\text{int}}\bigr{)}(z)\,=\,x+L(x^{\star})$
is another projection from $\mathbb{R}^{4}$ to $\mathbb{R}^{2}$. In fact, every projection from the total space $\mathbb{R}^{4}$ to direct space $\mathbb{R}^{2}$ can be written in this way. Changing projections is exactly the same as adding a linear map from internal space to direct space, which is a standard situation in the treatment of deformed model sets; see [4, Ex. 9.9] and references given there.
The coordinates of $x^{\star}$ are weakly pattern-equivariant functions on our Delone set. If two points have identical neighborhoods, their star coordinates are close but not identical. However, the difference $x_{1}^{\star}-x_{2}^{\star}$, where $x_{1}$ and $x_{2}$ are nearby points in our Delone set, is strongly pattern-equivariant. This means that the coboundaries of the coordinates of $x^{\star}$ (viewed as $1$-cochains on the tiling obtained by connecting the vertices of our Delone set) are linearly independent, asymptotically negligible classes in $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R})$. The span of the linear functions from internal space to direct space is then a $4$-dimensional space of asymptotically negligible classes in $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$. However, $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$ is only $4$-dimensional, so all asymptotically negligible classes are obtained in this way.
By reprojecting the CAP Delone set in arbitrary ways, we obtain Delone sets whose shape class differ from the CAP shape class by arbitrary asymptotically negligible elements of $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$. This means that every tiling that is topologically conjugate to the CAP Delone set is MLD to such a reprojection of the CAP Delone set, as claimed in Theorem 3.
Figure 10. Hat tiling with superimposed meta-tiles and control points. The latter are obtained by reprojection of the control points of the CAP tiling.
Remark 15.
Under topological conjugacy, singular points of the MEF must be mapped onto each other, preserving the cardinality of the fibres. The set of points in the MEF having a singular fibre (with more than one element) is given by $S=(\mathbb{R}^{2}\times\partial W)/\mathcal{L}$, where $\partial W$ is the set of all boundary points of the window (including internal boundaries), and $\mathbb{R}^{2}$ are the translations in the direction of direct space. This is a subset of measure zero of the MEF $(\mathbb{R}^{2}\times\mathbb{R}^{2})/\mathcal{L}$. Due to the fractal boundaries, a reflection preserving the total window cannot also preserve the set $S$. This means that translations of the hull in direct space direction cannot commute with a reflection, for any direction of projection, even if they do so for the MEF itself. Hence, a reflection can preserve the spectrum under certain conditions, but it cannot be a topological conjugacy. $\Diamond$
As an application, we determine the reprojected control points for the Hat tiling. Figure 10 shows a Hat tiling with superimposed meta-tiles, along with their control points. These control points are determined as follows. Let $\{t_{1},t_{2},t_{3},t_{4}\}$ be a basis of the return module $R$ of the CAP tiling, consisting of actual return vectors. For each $t_{i}$, one then finds a patch in a CAP tiling containing two translation-equivalent tiles which are a vector $t_{i}$ apart. There is then an equivalent patch in the Hat meta-tile tiling, where the two equivalent tiles are a vector $t_{i}^{\hskip 0.5pt\prime}$ apart. For instance, on the right of Figure 10, there are three hexagon meta-tiles with an orange edge at their lower right. Their distances, together with the corresponding preimages in the CAP tiling, could be used as pairs $(t_{i},t_{i}^{\hskip 0.5pt\prime})$ (complemented with rotated such pairs). The map $t_{i}\mapsto t_{i}^{\hskip 0.5pt\prime}$ can now be extended linearly to a map from the return module of the CAP tiling (of rank $4$) to the return module of the Hat tiling (of rank $2$). This map is the reprojection map for the control points. Each control point of a CAP tiling is mapped to a unique control point of the Hat tiling. From the (colored) control point set of the Hat tiling, the Hat tiling itself can be locally reconstructed. Consequently, the Hat tiling and the colored Delone set with the reprojected control points are MLD. Note that if the control points are distributed over several orbits under the return module, the reprojection procedure becomes considerably more involved.
4. Comparing different shapes
In this section, we compare several versions of the Hat tiling, in particular the CAP tiling and the four named tilings from [17] whose meta-tile vertices live on a lattice: the Chevron, the Comet, the Turtle and the original Hat. We will compute the expansions and rotations needed to make these versions topologically conjugate and see how their underlying lattices are rotated relative to one another. We will compute the return module for each one and see how its dual sits inside the already computed spectrum of the CAP tiling. Note that we work here with the coordinates of the original Hat and meta-tiles, before rescaling the return module as for the computation of the window in Figure 8.
Finally, we consider arbitrary elements of the Hat family considered in [17], with all tile edges either being vertical or horizontal, up to rotation by multiples of $60$ degrees. For each such shape, the spectrum for the resulting tiling by meta-tiles is rotated relative to the coordinate axes. The spectrum for the other way to tile the plane, using a hierarchical system of anti-meta-tiles, is rotated in the opposite direction by the same angle. As long as the angle of rotation is not a multiple of $30$ degrees, the two rotated patterns are different. That is, it is usually possible to distinguish between the two LI classes of possible tilings by spectral means. This is the case for all but two members of the Hat family: the Turtle and a shape that has not yet been named.
The geometry of the Hat tile is shown in [17, Fig. 3.2]. In the basic shape, the orange edges have length $\sqrt{3}/2$, while the black edges have length $1/2$. More generally, we can let the orange and black edges involve displacements by complex numbers $\alpha$ and $\beta$. The Hat is then a polygon whose displacements are
$\alpha,\,\xi\alpha,\,-\hskip 0.5pt\mathrm{i}\xi\beta,\,-\hskip 0.5pt\mathrm{i}\xi^{2}\beta,\,-\xi^{2}\alpha,\,-\xi\alpha,\,-\hskip 0.5pt\mathrm{i}\xi\beta,\,-\hskip 0.5pt\mathrm{i}\beta,\,-\alpha,\,\xi^{2}\alpha,\,\hskip 0.5pt\mathrm{i}\xi^{2}\beta,\,2\hskip 0.5pt\mathrm{i}\xi\beta,\,\text{and }\hskip 0.5pt\mathrm{i}\beta\hskip 0.5pt.$
Varying $\alpha$ and $\beta$ gives our (real) $4$-parameter family of shape changes that respect rotational symmetry.
In order for an anti-Hat to fit next to a Hat, the edges of the anti-Hat must also involve $\alpha$ and $\beta$. However, the reflection of the Hat about the horizontal axis has edge vectors $\bar{\alpha}$ and $-\hskip 0.5pt\mathrm{i}\bar{\beta}$ times powers of $\xi$. That is, the anti-Hat can only be the reflection of the Hat about a horizontal axis if $\alpha$ and $\beta$ are real. This is why the authors of [17] only considered real ratios $\alpha:\beta$. We will eventually consider arbitrary ratios $\alpha:\beta$, but for now we consider four named versions in which the vertices of the meta-tiles lie on the triangular lattice $\hskip 0.5pt\mathbb{Z}[\xi]$, and the self-similar CAP tiling as well as Socolar’s version from [19].
  • When $\alpha=1$ and $\beta=0$, the Hat degenerates to a Chevron, which is the union of four equilateral triangles with a common vertex. After rotating by $60$ degrees (which is merely changing which orientation is considered standard), the shape $(A,B,X,F,L)$ is given by the row vector
    $(1+\xi,1+\xi,1,1,0)\,=\,\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2}\hskip 0.5pt.$
  • The original Hat has $\alpha=\sqrt{3}/2$ and $\beta=1/2$. After rotating by $90$ degrees, our shape vector is
    $(3,3,1,1,1)\,=\,\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2}+\boldsymbol{\mu}_{3}\,\sim\,\frac{\raisebox{-2.0pt}{$3$}}{\raisebox{0.5pt}{$2$}}\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2}\hskip 0.5pt,$
    where $x\sim y$ means that $x$ and $y$ are cohomologous.
  • The Turtle has $\alpha=1/2$ and $\beta=\sqrt{3}/2$. After rotating by $120$ degrees, our shape vector is
    $\bigl{(}2(1+\xi^{5}),2(1+\xi^{5}),1,1,1+\xi^{5}\bigr{)}\,=\,\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2}+(1+\xi^{5})\boldsymbol{\mu}_{3}\,\sim\,\frac{\raisebox{-2.0pt}{$3+\xi^{5}$}}{\raisebox{0.5pt}{$2$}}\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2}\hskip 0.5pt.$
  • When $\alpha=0$ and $\beta=1$, the Hat degenerates to a Comet, which is the union of eight equilateral triangles. After rotating by $90$ degrees, $(A,B,X,F,L)$ equals
    $(\xi,\xi,\xi,\xi,2)\,=\,\xi(\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2})+2\boldsymbol{\mu}_{3}\,\sim\,(\xi+1)\boldsymbol{\mu}_{1}+\xi\boldsymbol{\mu}_{2}\hskip 0.5pt.$
  • As noted earlier, the CAP tiling has shape vector
    $(3\phi,3\phi,1,\phi+\xi,2\phi)\,=\,\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}+2\phi\boldsymbol{\mu}_{3}\,\sim\,(1+\phi)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}\hskip 0.5pt.$
  • We note in passing that the shape change used to describe Socolar’s golden Key tiles [19] is described by the shape vector
    $\begin{split}(3\phi+2-\xi,3\phi+2-\xi,0.\phi+\xi,2\phi+2)\,&=\,(\phi+\xi)\boldsymbol{\mu}_{2}+(2\phi+2)\boldsymbol{\mu}_{3}\\ &\sim\,(1+\phi)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}\hskip 0.5pt,\end{split}$
    which is thus cohomologous to that of the CAP tiling. This implies that the golden Key tiling is MLD to the CAP tiling.
Let us next compute the return modules for each tiling. The return module is generated by collections of edges that correspond to loops in the AP complex. That is, we want chains that are in the kernel of the boundary map $\partial_{1}$. A basis for this kernel is
  1. (1)
    $\;A+r^{3}B$ (times powers of $r$).
  2. (2)
    $\;A+(r^{3}-r^{5})X$ (times powers of $r$).
  3. (3)
    $\;(1-r^{2})F$ (times powers of $r$), and
  4. (4)
    $\;(r+1)\bigl{(}(1-r^{3})X+L\bigr{)}$ (times powers of $r$).
However, we must have
$A\,=\,B\,=\,(1+\xi^{5})X+(\xi+\xi^{2})F+(1+\xi^{5})L\hskip 0.5pt.$
With this, our return module is generated by
$(\xi+\xi^{2})F+(1+\xi^{5})L,\quad(1-\xi^{2})F,\;\hbox{ and }\,(\xi+1)(2X+L)\hskip 0.5pt,$
times powers of $\xi$. By taking linear combinations, this reduces to
$(\xi+1)F,\quad(\xi+1)L,\;\hbox{ and }\,2(\xi+1)X.$
  • In the Chevron, Hat, Turtle and Comet, $X=F$ is a power of $\xi$, while $L$ is in $\hskip 0.5pt\mathbb{Z}[\xi]$. This makes the return module $(1+\xi)\hskip 0.5pt\mathbb{Z}[\xi]$. These are the only members of the Hat family of tilings where the return module has rank 2.
  • In the CAP tiling, the generators are $(1+\xi)$ times 2, $\phi+\xi$ and $2\phi$. We could just as well take the generators to be $(1+\xi)$ times 2, $\phi-\xi$ and $2\phi$. The resulting module is invariant under multiplication by $\xi$ and also by $\phi$, the latter because
[…]
    which is to say that it is a $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$-module. However, over $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, the numbers $2$ and $2\phi$ are multiples of $\phi-\xi$, since
[…]
    Thus, our module is $(1+\xi)(\phi-\xi)\hskip 0.5pt\mathbb{Z}[\xi,\phi]$. Since $\phi^{2}(1+\xi)(\phi-\xi)=3\phi+2-\xi=t_{0}$, see Eq. (15), and $\phi^{2}$ is a unit of $\hskip 0.5pt\mathbb{Z}[\xi,\phi]$, this agrees with our findings in Section 3.
Next, we determine the rescalings needed to make these tilings conjugate. We wish to decompose the cohomology class $S$ of each shape vector in terms of the left eigenvectors $\ell_{\lambda}=(\lambda-\xi)\boldsymbol{\mu}_{1}+2\boldsymbol{\mu}_{2}$ of the substitution operator as in equation (14), where $\lambda=\phi^{\pm 2}$. We get the coefficient of $\ell_{\lambda}$ by taking an inner product of $S$ with the right-eigenvector $r_{\lambda}$ and normalizing, thus obtaining the coefficient of $\ell_{\lambda}$ as
$\frac{S\cdot r_{\lambda}}{\ell_{\lambda}\cdot r_{\lambda}}\,=\,\frac{S\cdot r_{\lambda}}{(3-2\xi)\lambda+(3\xi-2)}\hskip 0.5pt.$
The values of the inner product $S\cdot r_{\lambda}$ for $\lambda=\phi^{2}$ and $\lambda=\phi^{-2}$ are shown in Table 3.
Table 3. A comparison of the tilings discussed in this paper. The second and third columns are the coefficients of $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, while the last two columns show the inner product of the shape class $S$ and the right eigenvectors $r_{\lambda}$ from (14).
Tiling $\boldsymbol{\mu}_{1}$ $\boldsymbol{\mu}_{2}$ $S\cdot r_{\phi^{2}}$ $S\cdot r_{\phi^{-2}}$
Chevron 1 1 $\phi^{2}$ $\phi^{-2}$
Hat 3/2 1 $-\frac{1}{2}\xi\phi^{2}(\phi-\xi^{5})^{3}$ $\frac{1}{2}\xi\phi^{-2}(\phi-\xi)^{3}$
Turtle $(3+\xi^{5})/2$ 1 $\frac{\sqrt{5}}{2}\hskip 0.5pt\xi^{5}\phi^{2}(\phi-\xi^{5})$ $\frac{\sqrt{5}}{2}\hskip 0.5pt\xi^{5}\phi^{-2}(\phi-\xi)$
Comet $1+\xi^{5}$ 1 $\phi^{2}(\phi-\xi)$ $-\phi^{-2}(\phi-\xi^{5})$
CAP $\phi+1$ $\phi+\xi$ $\sqrt{5}\,\phi^{2}$ 0
From here on, we take the CAP tiling as our reference. To make the other tilings topologically conjugate to it, we must multiply the shape vectors of the Chevron, Hat, Turtle, and Comet by $\sqrt{5}$, $\xi^{2}\sqrt{5}\hskip 0.5pt(\phi-\xi)^{3}/4$, $\xi(\phi-\xi)$, and $\sqrt{5}\hskip 0.5pt(\phi-\xi^{5})/2$, respectively. Note that the argument of $\phi-\xi^{5}$ is $\arctan\bigl{(}\sqrt{3/5\hskip 0.5pt}\,\bigr{)}$, which is not a rational angle. With these rotations (and rescalings), the underlying lattices of the Comet, Hat, Turtle, and Chevron have become incommensurate. What connects them is the higher-rank return module of the CAP tiling.
Let $L_{0}=\hskip 0.5pt\mathbb{Z}[\xi]$ be the standard triangular lattice, so $L_{0}^{*}=\frac{2\hskip 0.5pt\hskip 0.5pt\mathrm{i}}{\sqrt{3}}L_{0}$ is its dual. The spectrum of the CAP tiling is the dual of the return module
$(1+\xi)(\phi-\xi)(L_{0}\oplus\phi L_{0})\,=\,(1+\xi)(\phi-\xi)\hskip 0.5pt\mathbb{Z}[\xi,\phi]\hskip 0.5pt,$
and is therefore
$\frac{(1+\xi^{5})^{-1}(\phi-\xi^{5})^{-1}}{\sqrt{5}}(L_{0}^{*}\oplus\phi L_{0}^{*})\,=\,\frac{(1+\xi)(\phi-\xi)}{6\sqrt{5}}\bigl{(}L_{0}^{*}\oplus\phi L_{0}^{*}\bigr{)}.$
For the compuation of the window (Figure 8), we have divided by the prefactor $(1+\xi)(\phi-\xi)$ (and the unit $\phi^{2}$).
Of course, the duals to the return modules of the Chevron, Hat, Turtle and Comet tilings must be in the spectrum of those tiling spaces, and so must be in the spectrum of the CAP tiling. We check that this is indeed the case, despite the incommensurate nature of those return modules.
  • The return module of the rescaled Chevron is $\sqrt{5}(1+\xi)L_{0}$, whose dual is $\frac{1}{3\sqrt{5}}(1+\xi)L_{0}^{*}$.
  • The return module of the rescaled Hat is $\frac{\sqrt{5}}{4}(1+\xi)(\phi-\xi)^{3}L_{0}$, whose dual is
    $\frac{1}{6\sqrt{5}}(1+\xi)(\phi-\xi)^{3}L_{0}^{*}$.
  • The return module of the rescaled Turtle is $(1+\xi)(\phi-\xi)L_{0}$, whose dual is
    $\frac{1}{6}(1+\xi)(\phi-\xi)L_{0}^{*}$.
  • The return module of the rescaled Comet is $\frac{\sqrt{5}}{2}(1+\xi)(\phi-\xi^{5})L_{0}$, whose dual is
    $\frac{1}{3\sqrt{5}}(1+\xi)(\phi-\xi^{5})L_{0}^{*}$.
To see that these are all contained in $\frac{1}{6\sqrt{5}}(1+\xi)(\phi-\xi)(L_{0}^{*}\oplus\phi L_{0}^{*})$, we note that
(21) $\begin{split}\frac{1+\xi}{3\sqrt{5}}&\,=\,(\phi-\xi^{5})\,\frac{(1+\xi)(\phi-\xi)}{6\sqrt{5}}\hskip 0.5pt,\\ \frac{(1+\xi)(\phi-\xi)^{3}}{6\sqrt{5}}&\,=\,(\phi-\xi)^{2}\,\frac{(1+\xi)(\phi-\xi)}{6\sqrt{5}}\hskip 0.5pt,\\ \frac{(1+\xi)(\phi-\xi^{5})}{6}&\,=\,(2\phi-1)\,\frac{(1+\xi)(\phi-\xi)}{6\sqrt{5}}\hskip 0.5pt,\\ \frac{(1+\xi)(\phi-\xi^{5})}{3\sqrt{5}}&\,=\,(\phi-\xi^{5})^{2}\,\frac{(1+\xi)(\phi-\xi)}{6\sqrt{5}}\hskip 0.5pt.\end{split}$
Finally, we consider the spectra of the entire family of Hat-like tilings considered in [17] and prove Theorem 6. For this purpose, we express everything in terms of the (real) parameters $\alpha$ and $\beta$. In order to consider all of these models on the same footing, we declare the standard orientation of the tile to be one where the edges of length $\alpha$ and $\beta$ are horizontal (resp. vertical), up to rotation by multiples of $60$ degrees. (For some of the named tilings, this disagrees with our previous conventions.) With this choice, the displacements of the five basic edges $A$, $B$, $X$, $F$ and $L$ in terms of $\alpha$ and $\beta$ are
$A\,=\,B$ $\,=\,(\xi+\xi^{2})\alpha+3\hskip 0.5pt\hskip 0.5pt\mathrm{i}\beta\hskip 0.5pt,$
(22) $X\,=\,F$ $\,=\,\xi(\alpha+\hskip 0.5pt\mathrm{i}\beta)\hskip 0.5pt,$
$L$ = $\,=\,2\hskip 0.5pt\hskip 0.5pt\mathrm{i}\beta\hskip 0.5pt.$
Using Eq. (12), our shape class is then
(23) $\begin{split}S\,&=\,\alpha(\xi+\xi^{2},\xi+\xi^{2},\xi,\xi,0)+\hskip 0.5pt\mathrm{i}\beta(3,3,\xi,\xi,2)\\[2.84526pt] &=\,(\alpha+\hskip 0.5pt\mathrm{i}\beta)(\xi\boldsymbol{\mu}_{1}+\xi\boldsymbol{\mu}_{2})+2\hskip 0.5pt\hskip 0.5pt\mathrm{i}\beta\boldsymbol{\mu}_{3},\end{split}$
which is cohomologous to $\xi(\alpha+\hskip 0.5pt\mathrm{i}\beta)(\boldsymbol{\mu}_{1}+\boldsymbol{\mu}_{2})+\hskip 0.5pt\mathrm{i}\beta\boldsymbol{\mu}_{1}$. Note that we can achieve any linear combination of $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$ by adjusting $\alpha$ and $\beta$ appropriately. This shows that all shape changes that preserve rotational symmetry are MLD to changes in $\alpha$ and $\beta$.
We now compute how the spectrum of this tiling aligns with the coordinate axes. Taking the inner product of the shape class $S$ with $r_{\phi^{2}}$ from (14) gives
$\xi\phi^{2}(\alpha+i\beta)+\hskip 0.5pt\mathrm{i}\beta(\phi^{2}-\xi)\,=\,\xi\phi^{2}\bigl{(}\alpha+\hskip 0.5pt\mathrm{i}\beta(\phi-\xi)\bigr{)}.$
That is, the $(\alpha,\beta)$ tiling is topologically conjugate to a rescaling of the CAP tiling by the factor $\xi\bigl{(}\alpha+\hskip 0.5pt\mathrm{i}\beta(\phi-\xi)\bigr{)}/\sqrt{5}$. The factors of $\xi$ and $\sqrt{5}$ are irrelevant, as the tilings are invariant under rotation by $60$ degrees and we are keeping track of angles, not scale.
The spectrum of the CAP tiling is itself rotated by the argument of $(\phi-\xi)$ relative to the axes of the CAP tiling, [^4] The factor of $(1+\xi)=\sqrt{3}\hskip 0.5pt\mathrm{i}\hskip 0.5pt\xi^{5}$ cancels the factor of $\hskip 0.5pt\mathrm{i}$ in $L_{0}^{*}=\frac{2\hskip 0.5pt\mathrm{i}}{\sqrt{3}}L_{0}$. so the spectrum of the $(\alpha,\beta)$ tiling is rotated by the argument of $\bigl{(}\alpha+\hskip 0.5pt\mathrm{i}\beta(\phi-\xi)\bigr{)}(\phi-\xi)$ relative to the original $x$ axis. A simple computation shows that
(24) $\bigl{(}\alpha+\hskip 0.5pt\mathrm{i}\beta(\phi-\xi)\bigr{)}(\phi-\xi)\,=\,\frac{\raisebox{-2.0pt}{$1$}}{\raisebox{0.5pt}{$2$}}\bigl{(}\sqrt{5}(\alpha+\sqrt{3}\beta)+\hskip 0.5pt\mathrm{i}(\beta-\sqrt{3}\alpha)\bigr{)}.$
For the argument of this number to be a multiple of $30$ degrees, the ratio
$\frac{\beta-\sqrt{3}\alpha}{\sqrt{5}(\alpha+\sqrt{3}\beta)}$
must be 00, ±$\pm 1/\sqrt{3}$, or ±$\pm\sqrt{3}$.
Since both $\alpha$ and $\beta$ are non-negative, the ratio is always between $-\sqrt{3/5}$ and $1/\sqrt{15}$, and so can never equal $\sqrt{1/3}$ or ±$\pm\sqrt{3}$. That leaves $-1/\sqrt{3}$, which is achieved when
$\beta\,=\,\frac{\sqrt{5}-2}{\sqrt{3}}\alpha\hskip 0.5pt,$
and 00, which is achieved when
$\beta\,=\,\sqrt{3}\hskip 0.5pt\alpha\hskip 0.5pt.$
The shape corresponding to the first ratio is similar to the Chevron, since $(\sqrt{5}-2)/\sqrt{3}\approx 0.136$ is small. This shape does not (yet) have a name. The shape corresponding to the second ratio is the Turtle. Although the two LI classes for the Turtle have the same spectrum, the argument of Remark 15 shows that they are not topologically conjugate, nor are the two LI classes of the other shape.
This completes the proof of Theorem 6.
5. Some comments and questions
Unlike previous examples, which often were limit-periodic, the CAP tiling is quasiperiodic in the sense of mean almost-periodic measures, which have recently been analyzed by Lenz, Spindeler and Strungaru [13]. As such, its pure-point diffraction corresponds to the mean almost periodicity of the (possibly color-weighted) Dirac comb on the control points. This quasiperiodic nature is preserved for the reprojected point sets, and also the Hat tiling is thus quasiperiodic in this sense.
Consequently, the diffraction measure of the Hat tiling is also pure point. Since the control points live on a lattice, the corresponding diffraction measure is the lattice-periodic repetition of a finite motif (or block), with the dual lattice as lattice of periods, by an application of [4, Thm. 10.3]. This is an interesting feature in the light of the intrinsic aperiodicity, though well known from other examples such as the aperiodic Fibonacci chain with two types of weights on $\hskip 0.5pt\mathbb{Z}$, which then has a periodic diffraction with dense peaks.
For the Hat, the motif can be represented as a finite point measure on a fundamental domain of the dual lattice that consists of a dense set of Dirac peaks. They then encode the modulation due to the quasiperiodic nature of the tiling, as also analyzed numerically in [19]. Some more detailed analysis of this feature seems of interest.
The quasiperiodic nature of the Hat tiling with its Euclidean internal space leads to a deformation structure that differs fundamentally from that of a limit-periodic system with its $p$-adic (or related) internal space. Specifically, having a Euclidean internal space allows for reprojections. Reprojections that preserve rotational symmetry then allow for an infinite family of monotiles.
We close with several open questions concerning the development of monotiles.
  • Do there exist other monotiles, not conjugate to the Hat, whose quasiperiodicity similarly implies the existence of an infinite family of monotiles?
  • Does there exist a disk-like monotile for which the tilings by isometric copies of the tile define a unique and aperiodic LI class?
  • Does there exist an aperiodic disk-like chiral monotile whose tilings only involve rotations and translations (but not reflections) of the basic tile?
  • Does there exist a disk-like chiral monotile for which the tilings by rotations and translations (but not reflection) of the tile define a unique and aperiodic LI class?
Note: Within a month of the initial release of this paper, Smith et al [18] answered the third question to the affirmative. They constructed a new aperiodic monotile, called the Spectre, whose tilings use $12$ rotations of the basic tile but no reflections. Tilings built from these tiles form two LI classes, each $6$-fold rotationally symmetric, but rotated against each other by $122\pi/$ 12. The Spectre tiling is quasiperiodic (in the sense we used above) and admits shape conjugacies that respect $6$-fold rotational symmetry. However, these deformations are not $12$-fold rotationally symmetric and do not yield an infinite family of monotiles.
Acknowledgements
It is our pleasure to thank Dirk Frettlöh, Craig Kaplan and Jan Mazáč for discussions. This work was supported by the German Research Council (Deutsche Forschungsgemeinschaft, DFG) under SFB-TRR 358/1 (2023) – 491392403.